# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 8.1. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$
(8.3) $0\to\Sigma^{2k,k}B_{0}^{\mathbb{R}}(\tfrac{k}{2})\to B_{0}^{\mathbb{R}}(k)\to B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1)\otimes(\mathcal{A}(1)_{\mathbb{R}}\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{R}}^{\vee}\to 0$
Applying $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(-)$ gives a long exact sequence of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules. Moreover, after killing $v_{1}$-torsion the connecting homomorphism is trivial, giving a short exact sequence whose middle term is $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$. Therefore, the Ext group in question decomposes into the Ext groups of the kernel and cokernel of the short exact sequence (8.3). By the inductive hypothesis, we know that
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(\Sigma^{2k,k}B_{0}^{\mathbb{R}}(\tfrac{k}{2})\right)\cong\Sigma^{2k,k}\left(\Sigma^{2k-4,k-2}Z_{\mathbb{\alpha}(k/2)}^{\mathbb{R}}\oplus\bigoplus_{j=0}^{2k-8}(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H)\right).$
@@
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))^{\vee}_{\mathbb{R}}\right)\cong\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1)\right).$
The group $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1))$ was calculated in Lemma 8.1. Thus the right-hand side and left-hand side assemble to give
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))\cong\Sigma^{4k-4,2k,2}Z_{\alpha(k)}^{\mathbb{R}}$ $\oplus\bigoplus_{j=0}^{2k-8}\left(\Sigma^{4j+2k,2j+2k}H\oplus\Sigma^{4j+4k,2j-2+2k}H\right)$
@@
using that $\alpha(\frac{k}{2})=\alpha(k)$. Notice that for degree reasons, any $\rho$-divisible class coming from $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1))$ is $v_{1}$-torsion. Combining and reindexing the $H$-towers in the second and third summands gives the result.
@@
(8.4) $0\to\Sigma^{2(k-1),k-1}B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\to B_{0}^{\mathbb{R}}(k)\to B_{1}^{\mathbb{R}}(\tfrac{k-1}{2}-1)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{R}}^{\vee}\to 0$
Again, modulo $v_{1}$-torsion the connecting homomorphism in the long exact sequence obtained by applying $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(-)$ is trivial, giving a short exact sequence whose middle term is $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k)).$ We proceed by analyzing the Ext groups of the kernel and cokernel of the short exact sequence (8.4). A change of rings isomorphism gives the right-hand side:
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k-1}{2}-1)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))^{\vee}_{\mathbb{R}}\right)\cong\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k-1}{2}-1)\right).$
The group $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k-1}{2}-1)\right)$ was calculated in Lemma 8.1. To calculate the left-hand side, we can use the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\right)$. This spectral sequence is obtained by applying the functor $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes-\right)$ to the cellular filtration for $B_{0}^{\mathbb{R}}(1)$ and has signature
$E_{1}=\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\right)\otimes\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\bar{\xi}_{1}],[\bar{\tau}_{1}]\}\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\right).$
@@
(8.5) $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\right)\cong\Sigma^{2k-6,k-3}Z_{\alpha(k)-1}^{\mathbb{R}}\oplus\bigoplus_{j=0}^{2k-10}\left(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H\right),$
using that $\alpha(\tfrac{k-1}{2})=\alpha(k)-1$. This splitting descends to a splitting of the spectral sequence, and we may analyze each summand of (8.5) individually. The left-hand summand is handled by 7.10, as it is isomorphic to the $\textbf{aAHSS}\left(\Sigma^{2k-6,k-3}B_{0}^{\mathbb{R}}(1)^{\otimes\alpha(k)}\right)$, so we have a summand of $\Sigma^{2k-2,k-1}Z_{\alpha(k)}^{\mathbb{R}}$ in $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\right)$. The algebraic Atiyah–Hirzebruch spectral sequence on the right-hand summand collases on $E_{2}$. The remaining classes in Adams filtration 0 are $v_{1}$-torsion, and so the result is isomorphic to the original right-hand summand. Thus, we have an isomorphism:
$\Sigma^{2(k-1),k-1}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\right)$
@@
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))\cong\Sigma^{4k-4,2k-2}Z^{\mathbb{R}}_{\alpha(k)}$ $\oplus\bigoplus_{j=0}^{2k-10}\left(\Sigma^{4j+2(k-1),2j+k-1}H\oplus\Sigma^{4j+2(k-1),2j-2+k-1}H\right)$
@@
As before, for degree reasons, any $\rho$-divisible class coming from $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}\left(\tfrac{k-1}{2}-1)\right)$ is $v_{1}$-torsion. Combining and reindexing the $H$-towers in the second and third summands gives the result. ∎
@@ 8.2. The ring of cooperations
In this subsection only, all results are implicitly computed modulo $v_{1}$-torsion. We assemble our results to compute the ring of cooperations $\pi_{**}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$. First, our computation in 8.2 gives the following:
@@
$E_{2}=\bigoplus_{k\geq 0}\textup{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\Sigma^{4k,2k}B_{0}^{\mathbb{R}}(k))\cong\bigoplus_{k\geq 0}\left(\Sigma^{4k-4,2k-2}Z_{\mathbb{\alpha}(k)}^{\mathbb{R}}\oplus\bigoplus_{j=0}^{4k-8}\left(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H\right)\right).$
@@
$-\otimes\mathbb{C}:\text{SH}(\mathbb{R})\to\text{SH}(\mathbb{C})$
[…]

### v2

@@ 8.1. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$
(8.1) $0\to\Sigma^{2k,k}B_{0}^{\mathbb{R}}(\tfrac{k}{2})\to B_{0}^{\mathbb{R}}(k)\to B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{R}}^{\vee}\to 0.$
Applying $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(-)$ gives a long exact sequence of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules. Moreover, after killing $v_{1}$-torsion the connecting homomorphism is trivial, giving a short exact sequence whose middle term is $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$. Therefore, the Ext group in question decomposes into the Ext groups of the kernel and cokernel of the short exact sequence (8.1). By the inductive hypothesis, we know that
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(\Sigma^{2k,k}B_{0}^{\mathbb{R}}(\tfrac{k}{2})\right)\cong\Sigma^{2k,k}\left(\Sigma^{2k-4,k-2}Z_{\mathbb{\alpha}(k/2)}^{\mathbb{R}}\oplus\bigoplus_{j=0}^{2k-8}(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H)\right).$
@@
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))^{\vee}_{\mathbb{R}}\right)\cong\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1)\right).$
The group $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1))$ was calculated in Lemma 8.1. Thus the right-hand side and left-hand side assemble to give
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))\cong\Sigma^{4k-4,2k,2}Z_{\alpha(k)}^{\mathbb{R}}$ $\oplus\bigoplus_{j=0}^{2k-8}\left(\Sigma^{4j+2k,2j+2k}H\oplus\Sigma^{4j+4k,2j-2+2k}H\right)$
@@
using that $\alpha(\frac{k}{2})=\alpha(k)$. Notice that for degree reasons, any $\rho$-divisible class coming from $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}(\tfrac{k}{2}-1))$ is $v_{1}$-torsion. Combining and reindexing the $H$-towers in the second and third summands gives the result.
@@
(8.2) $0\to\Sigma^{2(k-1),k-1}B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\to B_{0}^{\mathbb{R}}(k)\to B_{1}^{\mathbb{R}}(\tfrac{k-1}{2}-1)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{R}}^{\vee}\to 0.$
Again, modulo $v_{1}$-torsion the connecting homomorphism in the long exact sequence obtained by applying $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(-)$ is trivial, giving a short exact sequence whose middle term is $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k)).$ We proceed by analyzing the Ext groups of the kernel and cokernel of the short exact sequence (8.2). A change of rings isomorphism gives the right-hand side:
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k-1}{2}-1)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))^{\vee}_{\mathbb{R}}\right)\cong\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k-1}{2}-1)\right).$
The group $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}\left(B_{1}^{\mathbb{R}}(\tfrac{k-1}{2}-1)\right)$ was calculated in Lemma 8.1. To calculate the left-hand side, we can use the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\right)$. This spectral sequence is obtained by applying the functor $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes-\right)$ to (4.3) and has signature
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\right)\otimes\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\bar{\xi}_{1}],[\bar{\tau}_{1}]\}\implies\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\right).$
@@
(8.3) $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\right)\cong\Sigma^{2k-6,k-3}Z_{\alpha(k)-1}^{\mathbb{R}}\oplus\bigoplus_{j=0}^{2k-10}\left(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H\right),$
using that $\alpha(\tfrac{k-1}{2})=\alpha(k)-1$. This splitting descends to a splitting of the spectral sequence, and we may analyze each summand of (8.3) individually. The left-hand summand is handled by 7.6, as it is isomorphic to the $\textbf{aAHSS}\left(\Sigma^{2k-6,k-3}B_{0}^{\mathbb{R}}(1)^{\otimes\alpha(k)}\right)$, so we have a summand of $\Sigma^{2k-2,k-1}Z_{\alpha(k)}^{\mathbb{R}}$ in $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\right)$. The algebraic Atiyah–Hirzebruch spectral sequence on the right-hand summand collases on $\textup{E}_{2}$. The remaining classes in Adams filtration 0 are $v_{1}$-torsion, and so the result is isomorphic to the original right-hand summand. Thus, we have an isomorphism:
$\Sigma^{2(k-1),k-1}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(\tfrac{k-1}{2})\otimes B_{0}^{\mathbb{R}}(1)\right)$
@@
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))\cong\Sigma^{4k-4,2k-2}Z^{\mathbb{R}}_{\alpha(k)}$ $\oplus\bigoplus_{j=0}^{2k-10}\left(\Sigma^{4j+2(k-1),2j+k-1}H\oplus\Sigma^{4j+2(k-1),2j-2+k-1}H\right)$
@@
As before, for degree reasons, any $\rho$-divisible class coming from $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{R}}}(B_{1}^{\mathbb{R}}\left(\tfrac{k-1}{2}-1)\right)$ is $v_{1}$-torsion. Combining and reindexing the $H$-towers in the second and third summands gives the result. ∎
Remark 8.4.
The summand
$\bigoplus_{j=0}^{4k-8}\left(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H\right)$
appearing in this decomposition is a ghost of the failure to express these Ext groups in terms of Adams covers of kq and ksp. It was shown in [26] that the motivic Adams spectral sequence for $\text{H}\mathbb{Z}$ collapses at the $\textup{E}_{2}$-page and has signature
$\textup{E}_{2}^{s,f,w}=\text{Ext}^{s,f,w}_{\mathcal{A}(0)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})\cong\mathbb{F}_{2}[\rho,\tau^{2},h_{0}]/(\rho h_{0})\implies\pi_{s,w}^{\mathbb{R}}(\text{H}\mathbb{Z}).$
In the Adams covers for kq and ksp, one sees summands of $\pi_{**}^{\mathbb{R}}(\text{H}\mathbb{Z})$. In the context of our calculations, we only obtain summands of $\pi_{**}^{\mathbb{R}}(\text{H}\mathbb{Z}/\rho)$.
@@ 8.2. The ring of cooperations
In this subsection and the next, all results are implicitly computed modulo $v_{1}$-torsion. We assemble our results to compute the ring of cooperations $\pi_{**}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$. First, our computation in 8.2 gives the following:
@@
$\textup{E}_{2}^{s,f,w}=\bigoplus_{k\geq 0}\textup{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\Sigma^{4k,2k}B_{0}^{\mathbb{R}}(k))\cong\bigoplus_{k\geq 0}\left(\Sigma^{4k-4,2k-2}Z_{\mathbb{\alpha}(k)}^{\mathbb{R}}\oplus\bigoplus_{j=0}^{4k-8}\left(\Sigma^{4j,2j}H\oplus\Sigma^{4j,2j-2}H\right)\right).$
@@
$\otimes\mathbb{C}:\text{SH}(\mathbb{R})\to\text{SH}(\mathbb{C})$
which induces a highly structured morphism from the $\textbf{mASS}^{\mathbb{R}}(\text{kq}\otimes\text{kq})$ to the $\textbf{mASS}^{\mathbb{C}}(\text{kq}\otimes\text{kq})$. The latter spectral sequence collapses on the $\text{E}_{2}$-page [17, Corollary 3.43]. Complex Betti realization shows that there can be no differentials in the $\textbf{mASS}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$ with both the source and target being $\tau$-free. However, by Corollary 8.3 we see that every class on the $\textup{E}_{2}$-page is $\tau^{4}$-periodic. Hence there can be no differentials and the spectral sequence collapses. ∎
8.3. Application to the kq-resolution
We conclude by examining the $\text{E}_{1}$-page of the kq-resolution. Recall that this spectral sequence has signature
$\textup{E}_{1}^{s,f,w}=\pi_{s+f,w}^{\mathbb{R}}(\text{kq}\otimes\overline{\text{kq}}^{\otimes f})\implies\pi_{s,w}^{\mathbb{R}}(\mathbb{S}).$
We can determine each filtration line of the $\textup{E}_{1}$-page by an extension of the techniques used for the computation of $\pi_{**}^{\mathbb{R}}(\text{kq}\otimes\text{kq}).$ For each $n\geq 0$, the $\textbf{mASS}^{\mathbb{R}}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n})$ takes the form
$\text{E}^{s,f,w}_{2}=\text{Ext}^{s,f,w}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\text{H}_{**}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n}))\implies\pi_{s,w}^{\mathbb{R}}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n}).$
We begin by decomposing this $\textup{E}_{2}$-page.
Lemma 8.5.
There is a Künneth isomorphism
$\textup{H}_{**}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n})\cong\textup{H}_{**}(\textup{kq})\otimes\textup{H}_{**}(\overline{\textup{kq}})^{\otimes n}.$
Proof.
We induct on $n$. The Künneth spectral sequence [18] takes the form
$\text{E}_{2}=\text{Tor}^{\mathbb{M}_{2}^{\mathbb{R}}}(\textup{H}_{**}(\textup{kq}),\textup{H}_{**}(\overline{\textup{kq}}))\implies\textup{H}_{**}(\textup{kq}\otimes\overline{\textup{kq}}).$
Since $\text{H}_{**}(\text{kq})=(\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))^{\vee}$ is the $\mathbb{M}_{2}^{\mathbb{R}}$-linear dual of a finitely-generated free $\mathbb{M}_{2}^{\mathbb{R}}$-module, it is also free. Thus all higher Tor terms vanish, implying that the spectral sequence collapes.
Suppose the result is true for all $i<n$. We again have a Künneth spectral sequence which takes the form
$\text{E}_{2}=\text{Tor}^{\mathbb{M}_{2}^{\mathbb{R}}}(\textup{H}_{**}(\textup{kq}\otimes\overline{\text{kq}}^{\otimes n-1}),\textup{H}_{**}(\overline{\textup{kq}}))\implies\textup{H}_{**}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n}).$
Note that $\text{H}_{**}(\overline{\text{kq}})$ is also free over $\mathbb{M}_{2}^{\mathbb{R}}$, which one can see by the long exact sequence in homology associated to the defining cofiber sequence
$\mathbb{S}\to\text{kq}\to\overline{\text{kq}}.$
Thus the higher Tor terms vanish, implying that the spectral sequence collapses. This gives an isomorphism
$\text{H}_{**}(\text{kq}\otimes\overline{\text{kq}}^{\otimes{n-1}})\otimes\text{H}_{**}(\overline{\text{kq}})\cong\text{H}_{**}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n}).$
By induction, we have a Künneth isomorphism on the left hand factor, finishing the proof. ∎
Note that there is an isomorphism of $\mathcal{A}(1)^{\vee}$-comodules due to Proposition 2.10
(8.5) $\text{H}_{**}(\overline{\text{kq}})\cong\bigoplus_{k\geq 1}\Sigma^{4k,2k}B_{0}^{\mathbb{R}}(k).$
We can now deduce the $\textup{E}_{2}$-page of the $\textbf{mASS}^{\mathbb{R}}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n})$.
Proposition 8.6.
The $\textup{E}_{2}$-page of the $\textup{{mASS}}^{\mathbb{R}}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n})$ takes the form
$\textup{E}_{2}^{s,f,w}\cong\bigoplus_{K\in\mathcal{K}_{n}}\Sigma^{4|K|,2|K|}\textup{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(K)),$
where $\mathcal{K}_{n}=\{K=(k_{1},\dots,k_{n}):k_{j}\geq 1\textup{ for all }j\}$, $|K|=\sum_{j=1}^{n}k_{j}$, and $B_{0}^{\mathbb{R}}(K)=\bigotimes_{j=1}^{n}B_{0}^{\mathbb{R}}(k_{j}).$
Proof.
By Lemma 8.5 and the change of rings isomorphism, we may rewrite the $\textup{E}_{2}$-page as
$\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{R}}}^{***}(\text{H}_{**}(\text{kq})\otimes\text{H}_{**}(\overline{\text{kq}})^{\otimes n})\cong\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\text{H}_{**}(\overline{\text{kq}})^{\otimes n}).$
The identification of (8.5) allows us to rewrite the first factor of $\text{H}_{**}(\overline{\text{kq}})$ in terms of Brown–Gitler comodules, leaving us with
$\bigoplus_{k_{1}\geq 1}\Sigma^{4k_{1},2k_{1}}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{***}(B_{0}^{\mathbb{R}}(k_{1})\otimes\text{H}_{**}(\text{kq})^{\otimes n-1}).$
Rewriting the next factor of $\text{H}_{**}(\overline{\text{kq}})$ gives us
$\bigoplus_{k_{1}\geq 1}\Sigma^{4k_{1},2k_{1}}\left(\bigoplus_{k_{2}\geq 1}\Sigma^{4k_{2},2k_{2}}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{***}(B_{0}^{\mathbb{R}}(k_{1})\otimes B_{0}^{\mathbb{R}}(k_{2})\otimes\text{H}_{**}(\overline{\text{kq}})^{\otimes n-2})\right),$
which we may rewrite as
$\bigoplus_{k_{1},k_{2}\geq 1}\Sigma^{4(k_{1}+k_{2}),2(k_{1}+k_{2})}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{***}(B_{0}^{\mathbb{R}}(k_{1})\otimes B_{0}^{\mathbb{R}}(k_{2})\otimes\text{H}_{**}(\overline{\text{kq}})^{\otimes n-2}).$
The result follows by continuing in this manner and rewriting all factors of $\text{H}_{**}(\overline{\text{kq}})$ in terms of Brown–Gitler comodules. ∎
Note that we have described the $\text{E}_{2}$-page of the $\textbf{mASS}^{\mathbb{R}}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n})$ as a module over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, hence as a module over the $\text{E}_{2}$-page of the $\textbf{mASS}^{\mathbb{R}}(\text{kq})$.
Theorem 8.7.
The $\textup{{mASS}}^{\mathbb{R}}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n})$ collapses on the $\textup{E}_{2}$-page.
Proof.
Base change to $\mathbb{C}$ induces a highly structured morphism from the $\textbf{mASS}^{\mathbb{R}}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n})$ to the $\textbf{mASS}^{\mathbb{C}}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n})$. The latter spectral sequence collapses on the $\text{E}_{2}$-page [17, Section 4.2]. Complex Betti realization shows that there can be no differentials in the $\textbf{mASS}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$ with both the source and target being $\tau$-free. By the description of the $\text{E}_{2}$-page in Proposition 8.6 and the corresponding Ext groups in 8.2, we see that every class on the $\textup{E}_{2}$-page is $\tau^{4}$-periodic. Hence there can be no differentials and the spectral sequence collapses. ∎